# Learning the Hamiltonian Matrix of Large Atomic Systems

### v1

@@ Abstract
Graph neural networks (GNNs) have shown promise in learning the ground-state electronic properties of materials, subverting ab initio density functional theory (DFT) calculations when the underlying lattices can be represented as small and/or repeatable unit cells (i.e., molecules and periodic crystals). Realistic systems are, however, non-ideal and generally characterized by higher structural complexity. As such, they require large (10+ Å) unit cells and thousands of atoms to be accurately described. At these scales, DFT becomes computationally prohibitive, making GNNs especially attractive. In this work, we present a strictly local equivariant GNN capable of learning the electronic Hamiltonian (${\bm{H}}$) of realistically extended materials. It incorporates an augmented partitioning approach that enables training on arbitrarily large structures while preserving local atomic environments beyond boundaries. We demonstrate its capabilities by predicting the electronic Hamiltonian of various systems with up to 3,000 nodes (atoms), 500,000+ edges, $\sim$28 million orbital interactions (nonzero entries of ${\bm{H}}$), and $\leq$1error in the eigenvalue spectra. Our work expands the applicability of current electronic property prediction methods to some of the most challenging cases encountered in computational materials science, namely systems with disorder, interfaces, and defects.
@@ 1 Introduction
  • We develop a strictly local GNN-based architecture tailored for electronic property prediction from small to large scales. Our network leverages efficient SO(2)-convolutions and multi-headed attention, with learnable embeddings for every node and edge that are mapped to diagonal (${\bm{H}}_{i,i}$) and off-diagonal (${\bm{H}}_{i,j}$) blocks of the Hamiltonian. We provide the code for this implementation in [anonymized].
@@
We combine our network and augmented partitioning approach to treat a custom dataset of three large atomic structures in their amorphous phase. As amorphous materials encompass all aforementioned defect types, they are ideally suited to test and validate our model. In particular, we achieve 0.99-5.16 meV prediction accuracies on unseen samples, matching the ranges of previous studies for materials with orders of magnitude fewer atoms (Wang et al., 2024b). We also demonstrate how this error translates into downstream applications by matching the eigenvalues of ${\bm{H}}$ to within 0.55% relative L1 error on structures with 3,000 atoms, which require several (3.65) hours to compute using DFT. Our work advances applications of equivariant GNNs towards practical use cases in computational materials science.
@@ 2 Background & related work
The entries of the ground-state Hamiltonian matrix ${\bm{H}}$ are typically computed from first principles with DFT ((Kohn & Sham, 1965)). In several widely used codes, the wavefunctions are expanded into a basis $|\varphi\rangle$ of non-orthogonal atomic orbitals localized around atomic positions, each built, for example, from contracted Gaussian functions ((Kühne et al., 2020; Neese, 2011)). These orbitals transform like spherical harmonics under rotation $\hat{{\bm{r}}}\rightarrow\hat{{\bm{r}}}^{\prime}$: $Y^{l}_{m}(\hat{{\bm{r}}}^{\prime})=\sum_{m^{\prime}}{\bm{D}}^{l}_{mm^{\prime}}({\bm{R}})Y^{l}_{m^{\prime}}(\hat{{\bm{r}}})$. Here, $Y^{l}_{m}$ is the spherical harmonic of degree $l$ and order $m\in\{-l,\dots,l\}$. ${\bm{D}}^{l}_{mm^{\prime}}({\bm{R}})$ is the Wigner-D matrix of degree $l$ corresponding to the rotation ${\bm{R}}$, which transforms the corresponding spherical harmonic. $\hat{{\bm{r}}}$ and $\hat{{\bm{r}}}^{\prime}$ are normalized direction vectors.
@@ 2.2 Hamiltonian prediction models
$({\bm{x}}_{j}^{l_{1}}\otimes^{l_{3}}_{l_{1},l_{2}}Y^{l_{2}}(\hat{{\bm{r}}}_{ji}))^{l_{3}}_{m_{3}}=\sum_{m_{1},m_{2}}({\bm{x}}^{l_{1}}_{j})_{m_{1}}C^{l_{3},m_{3}}_{(l_{1},m_{1}),(l_{2},m_{2})}{h}_{l_{1},l_{2},l_{3}}Y^{l_{2}}_{m_{2}}(\hat{{\bm{r}}}_{ji}),$
where the $C^{l_{3},m_{3}}_{(l_{1},m_{1}),(l_{2},m_{2})}$ are the Clebsch-Gordan coefficients that are indexed by the order $m$ and degree $l$ of the input, filter, and output embeddings. The combination of feature ($x$) and geometric ($\hat{r}$) information along each edge encodes both the identity and structure of the system. These ‘Tensor Field Networks’ (TFNs) (Hodapp & Shapeev, 2023) achieve state-of-the-art accuracy on small molecule (Yu et al., 2023b) and crystalline (Gong et al., 2023) datasets. However, they are also computationally expensive. The network training scales with $\mathcal{O}(l_{max}^{6})$, where $l_{max}$ is the maximum degree of the angular momentum considered. As a consequence, E(3)-equivariant tensor product networks are difficult to apply beyond a few atoms (Zhang et al., 2024).
Recently, the computational cost of training equivariant GNNs has been significantly reduced by combining the benefits of data rotation and equivariant network operations. These approaches take advantage of the fact that when edges are rotated to align with a fixed axis ($y$ or $z$ depending on convention), the only non-zero spherical harmonic components are those of order $m$ = 0. By keeping track of the bond vectors and performing internal spherical rotations, complex SO(3) convolutions can thus be reduced to SO(2) linear convolution operations (Passaro & Zitnick, 2023). Furthermore, under these conditions, the Clebsch-Gordan coefficients exhibit a predictable sparsity pattern (non-zero only when $m_{3}$ = $\pm$ $m_{1}$). Altogether, the scaling reduces to $\mathcal{O}(l_{max}^{3})$ (Wang et al., 2024a), thus speeding up the network’s training, while enabling the use of higher-order angular momenta ($l_{max}$) and more parameters to capture finer, more complex details of the surrounding environment. An advanced SO(2) convolution network was developed by (Passaro & Zitnick, 2023) (eSCN), while an implementation of this approach on Hamiltonians by (Wang et al., 2024a) achieved better performance on custom crystalline 2D-material datasets compared to previous tensor field and invariant networks. The architecture was further expanded by (Liao et al., 2023) (EquiformerV2) with the inclusion of equivariant attention. However, the prediction of large disordered structures still remains an open problem in literature.
@@ 3.1 Orbital block processing
Figure 3: (a) Data transformation of the Hamiltonian matrix. Blocks of orbital interactions are first extracted from ${\bm{H}}$ and reshaped into input tensors which are transformed into the coupled basis. The tensor corresponding to each node and edge is expanded into a dimension $E_{n}/E_{e}$, and this set of initial embeddings for every node and edge is sent into (b) the node update block. The node features (${\bm{n}}_{X}$) are updated using a message-passing scheme. The edge features (${\bm{e}}_{X}$) are updated based on the learned ${\bm{n}}_{X}$. Z refers to atomic numbers, while rij is the set of scalar distances between atoms. (c) Illustration of how connectivity beyond the partition boundaries is incorporated through the addition of virtual edges ($e_{j^{\prime}\rightarrow i}$) from virtual nodes ($j^{\prime}$) to the labeled nodes. A set of edges from labeled and virtual nodes is shown for a single node ($i$) within the partition. Solid vertical lines indicate the partition boundaries and different colors represent different atomic species.
@@ 3.2 Network architecture
During the node update phase, each node $i$ receives messages from all its neighbors $j$, consisting of the concatenated embeddings ${\bm{n}}_{i}$, ${\bm{n}}_{j}$, and ${\bm{e}}_{ji}$. To enable fast tensor product operations in large atomic structures, we adopt eSCN convolutions (Passaro & Zitnick, 2023), with incoming messages rotated to align with the $z$-axis during these operations. Non-linearity is then introduced through a gate activation layer. On top of this, we also include dynamic multi-headed attention mechanisms, as in EquiformerV2 (Liao et al., 2023), which allows our network to learn better from highly dense local atomic environments. Finally, the resulting output messages are rotated back to their original orientations, aggregated onto the node $i$, and passed through a feedforward network to update its output embedding ${\bm{n}}_{i}$ corresponding to onsite (diagonal) Hamiltonian blocks.
To learn offsite (non-diagonal) blocks, we use concepts from the Hamiltonian prediction networks in (Gong et al., 2023),(Wang et al., 2024a) and introduce learnable embeddings for every edge, defined between pairs of atoms located within a distance rcut from each other. The updated node embeddings are used to update the edges through a similar process, without the attention layer. The predicted outputs are then post-processed back into the uncoupled basis with the reverse transformation: $T^{-1}({\bm{H}}_{\gamma\delta}^{ij})$ = ${\bm{H}}_{\alpha\beta}^{ij}$, such that they can be used to reconstruct the Hamiltonian matrix block-by-block.
@@ 3.3 Training on augmented partitions
The combination of large graphs, dense connectivity, and large tensor representations of nodes and edges, results in high memory consumption and long computation times per epoch during training (see Appendix H.2). If a distributed computing environment was used, the communication of intermediate values during the forward pass would incur significant overhead, impeding scalability (Wan et al., 2022). At the same time, graphs cannot be arbitrarily partitioned by removing inter-atomic connections. Doing so misinforms the network as it tries to fit the target data while aggregating inputs through an incorrect/incomplete graph structure. The inability to divide the graph into batches slows down the training process overall.
@@ 3.4 Dataset creation
To generate sufficiently rich training data, existing datasets typically sample molecules at various time steps of molecular dynamics (MD) trajectories (Yu et al., 2023a; Schütt et al., 2019; Christensen & lilienfeld, 2020) or generate multiple small perturbations of the atoms in a crystalline lattice (Li et al., 2022). Here, as a representative subset of realistic (disordered) materials, we consider amorphous crystals and take advantage of the fact that (1) every node has a different local atomic environment, and (2) each structure contains a large sampling of different motifs. A wide range of training data can thus be captured within a single sample.
@@ 4.1 Ablation studies of the training approach
We study the model’s ability to generalize to different configurations of large systems by predicting the Hamiltonian matrix ${\bm{H}}$ of a full a-HfO2 sample (structure 3 in Table 1), which remains unseen during the training process. For all subsequent HfO2 experiments, the augmented partitioning scheme is only applied during training on structure 2, while the ${\bm{H}}$ of the full unseen structure is predicted during inference (structure 3). We use an rcut of 8 Å. Errors are reported separately for nodes ($\epsilon_{n}$) and edges ($\epsilon_{e}$) to distinguish intra- and inter-atomic orbital interactions, which have very different magnitudes.
@@
$n^{\prime}\;\;\;\;\;n$ 5.18 1.66
$n^{\prime}\rightarrow n$ 2.29 0.20
@@
First, we demonstrate the improvement in accuracy resulting from the augmented partitioning approach introduced in Section 3.3. In particular, we examine the influence of virtual nodes/edges in Table 2. Compared to training with raw partitions, the addition of virtual nodes and edges reduces the node ($\epsilon_{node}$) and edge prediction error ($\epsilon_{edge}$) by $\sim$50and $\sim$88respectively. Such an improvement is expected, as raw partitions omit a large proportion of boundary edges and thus incorrectly capture atomic neighborhoods. We note that as $t_{slice}<r_{cut}$, the one-hop neighborhood connects all nodes within each partition.
Next, we explore the impact of augmented partitioning on accuracy by training on an increasingly sub-divided graph in Table 3. The total number of labeled atoms used for training remains the same (3,000), while the total number of labeled edges decreases with more partitions. Despite the different divisions ranging from 5 ($t_{slice}\simeq$12 Å) to 27 ($t_{slice}\simeq$2 Å) slices, $\mathbf{\epsilon_{n}}$ and $\mathbf{\epsilon_{e}}$ remain very close to the values obtained by training with the full graph ($t_{slice}$ = 52.346 Å). The prediction error is thus insensitive to partition size. Using multiple partitions also allows for effective batching, reducing the number of epochs required for convergence. For small slices, the reduced fraction of labeled connections along the $x$ direction does not affect the accuracy, as the remaining data along the $y$ and $z$ directions is sufficient to train the network.
@@ 4.2 Structural disorder
The strictly local architecture combined with augmented partitioning performs consistently across a diverse range of test datasets, with total errors ranging from 0.99 meV to 5.16 meV, for test structures containing 1000-3000 atoms and 200,000+ to 2,000,000+ edges. These values are comparable to what a previous study obtained (2.2 $meV$) using equivariant GNNs for smaller structures with $\leq$150 atoms per unit cell (Wang et al., 2024b). In all cases, the prediction errors also remain relatively insensitive to the partition size (Table 11 in Appendix I). In the case of PtGe, the use of small slices enables to push $r_{cut}$ to 16 Å  and remain within GPU memory limitations, further improving prediction accuracy. The error can be further reduced with more training data (Table 9 in Appendix I).
@@ 4.3 Compositional disorder
To demonstrate our network’s ability to treat such problems, we train a model on a single sub-stoichiometric a-HfO1.8 structure where 10% of the oxygen atoms were replaced by randomly distributed vacancies represented as ghost atoms. We then use it to predict unseen full structures with a stoichiometry of a-HfO1.7, a-HfO1.8, and a-HfO1.9. The $\epsilon_{n}$ and $\epsilon_{e}$ remain within a small range (2.48-2.60 $mE_{h}$ and 0.17-0.18 $mE_{h}$, respectively), regardless of the vacancy concentration and distribution, showing that the network generalizes very well to compositional disorder, despite being trained on only one configuration. Models trained on other vacancy concentrations perform similarly. Their results, along with details of the vacancy implementation, can be found in Appendix I.2.

### v2

# Learning the Electronic Hamiltonian of Large Atomic Structures

@@ Abstract
Graph neural networks (GNNs) have shown promise in learning the ground-state electronic properties of materials, subverting ab initio density functional theory (DFT) calculations when the underlying lattices can be represented as small and/or repeatable unit cells (i.e., molecules and periodic crystals). Realistic systems are, however, non-ideal and generally characterized by higher structural complexity. As such, they require large (10+ Å) unit cells and thousands of atoms to be accurately described. At these scales, DFT becomes computationally prohibitive, making GNNs especially attractive. In this work, we present a strictly local equivariant GNN capable of learning the electronic Hamiltonian (${\bm{H}}$) of realistically extended materials. It incorporates an augmented partitioning approach that enables training on arbitrarily large structures while preserving local atomic environments beyond boundaries. We demonstrate its capabilities by predicting the electronic Hamiltonian of various systems with up to 3,000 nodes (atoms), 500,000+ edges, $\sim$28 million orbital interactions (nonzero entries of ${\bm{H}}$), and $\leq$0.53error in the eigenvalue spectra. Our work expands the applicability of current electronic property prediction methods to some of the most challenging cases encountered in computational materials science, namely systems with disorder, interfaces, and defects.
Machine Learning, ICML
@@ 1 Introduction
  • We develop a strictly local GNN-based architecture tailored for electronic property prediction from small to large scales. Our network leverages efficient SO(2)-convolutions and multi-headed attention, with learnable embeddings for node and edges that are mapped to diagonal (${\bm{H}}_{i,i}$) and off-diagonal (${\bm{H}}_{i,j}$) blocks.
@@
We combine our network and augmented partitioning approach to treat a custom dataset of three large atomic structures in their amorphous phase. As amorphous materials encompass all aforementioned defect types, they are ideally suited to test and validate our model. In particular, we achieve 2.17-2.58 meV prediction accuracies on unseen samples, matching the ranges of previous studies for materials with orders of magnitude fewer atoms (Wang et al., 2024b). We also demonstrate how this error translates into downstream applications by matching the eigenvalues of ${\bm{H}}$ to within 0.53% relative L1 error on structures with 3,000 atoms, which require several (3.65) hours to compute using DFT. Our work advances applications of equivariant GNNs towards practical use cases in computational materials science.
@@ 2 Background & related work
The entries of the ground-state Hamiltonian matrix ${\bm{H}}$ are typically computed from first principles with DFT ((Kohn & Sham, 1965)). In several widely used codes, the wavefunctions are expanded into a basis $|\varphi\rangle$ of non-orthogonal atomic orbitals localized around atomic positions, each built, for example, from contracted Gaussian functions ((Kühne et al., 2020; Neese, 2011)). These orbitals transform like spherical harmonics under rotation $\hat{{\bm{r}}}\rightarrow\hat{{\bm{r}}}^{\prime}$: $Y^{l}_{m}(\hat{{\bm{r}}}^{\prime})=\sum_{m^{\prime}}{\bm{D}}^{l}_{mm^{\prime}}({\bm{R}})Y^{l}_{m^{\prime}}(\hat{{\bm{r}}})$. Here, $Y^{l}_{m}$ is the spherical harmonic of degree $l$ and order $m\in\{-l,\dots,l\}$. ${\bm{D}}^{l}_{mm^{\prime}}({\bm{R}})$ is the Wigner-D matrix of degree $l$ corresponding to the rotation ${\bm{R}}$, which transforms the spherical harmonic $\hat{{\bm{r}}}$ and $\hat{{\bm{r}}}^{\prime}$ are normalized direction vectors.
@@ 2.2 Hamiltonian prediction models
$({\bm{x}}_{j}^{l_{1}}\otimes^{l_{3}}_{l_{1},l_{2}}Y^{l_{2}}(\hat{{\bm{r}}}_{ji}))^{l_{3}}_{m_{3}}=\sum_{m_{1},m_{2}}C({\bm{x}}^{l_{1}}_{j})_{m_{1}}{h}_{l_{1},l_{2},l_{3}}Y^{l_{2}}_{m_{2}}(\hat{{\bm{r}}}_{ji}),$
where $C=C^{l_{3},m_{3}}_{(l_{1},m_{1}),(l_{2},m_{2})}$ are the Clebsch-Gordan coefficients that are indexed by the order $m$ and degree $l$ of the input, filter, and output embeddings. The combination of feature ($x$) and geometric ($\hat{r}$) information along each edge encodes both the identity and structure of the system. These ‘Tensor Field Networks’ (TFNs) (Thomas et al., 2018) achieve state-of-the-art accuracy on small molecule (Yu et al., 2023b) and crystalline (Gong et al., 2023) datasets. However, they are also computationally expensive. The network training scales with $\mathcal{O}(l_{max}^{6})$, where $l_{max}$ is the maximum degree of the angular momentum considered. As a consequence, E(3)-equivariant tensor product networks are difficult to apply beyond a few atoms (Zhang et al., 2024).
Recently, the cost of training equivariant GNNs has been significantly reduced by combining the benefits of data rotation and equivariant network operations. These approaches take advantage of the fact that when edges are rotated to align with a fixed axis ($y$ or $z$, depending on convention), the only non-zero spherical harmonic components are those of order $m$ = 0. By keeping track of the bond vectors and performing internal spherical rotations, complex SO(3) convolutions can thus be reduced to SO(2) linear convolution operations (Passaro & Zitnick, 2023). Furthermore, under these conditions, the Clebsch-Gordan coefficients exhibit a predictable sparsity pattern (non-zero only when $m_{3}$ = $\pm$ $m_{1}$). Altogether, the scaling reduces to $\mathcal{O}(l_{max}^{3})$ (Wang et al., 2024a), speeding up training, while allowing for higher-order angular momenta ($l_{max}$) and more parameters to capture finer, more complex details of the surrounding environment. An advanced SO(2) convolution network was developed by (Passaro & Zitnick, 2023) (eSCN), and was further expanded by (Liao et al., 2023) (EquiformerV2) with the inclusion of equivariant attention. An implementation of this architecture on Hamiltonians by (Wang et al., 2024a) achieved better performance on custom crystalline 2D-material datasets compared to previous tensor field and invariant networks. However, the prediction of large disordered structures still remains an open problem in literature.
@@ 3.1 Orbital block processing
Figure 3: (a) Data transformation of the Hamiltonian matrix. Blocks of orbital interactions are first extracted from ${\bm{H}}$ and reshaped into input tensors, which are transformed into the coupled basis. The tensor corresponding to each node and edge is expanded into a dimension $E_{n}/E_{e}$, and this set of initial embeddings for every node and edge is sent into (b) the node update block. The node features (${\bm{n}}_{X}$) are updated using a message-passing scheme. The edge features (${\bm{e}}_{X}$) are updated based on the learned ${\bm{n}}_{X}$. Z refers to atomic numbers, while rij is the set of scalar distances between atoms. (c) Illustration of how connectivity beyond the partition boundaries is incorporated through the addition of virtual edges ($e_{j^{\prime}\rightarrow i}$) from virtual nodes ($j^{\prime}$) to the labeled nodes. A set of edges from labeled and virtual nodes is shown for a single node ($i$) within the partition. Solid vertical lines indicate the partition boundaries, and different colors represent different atomic species.
@@ 3.2 Network architecture
During the node update phase, each node $i$ receives messages from all its neighbors $j$, consisting of the concatenated embeddings ${\bm{n}}_{i}$, ${\bm{n}}_{j}$, and ${\bm{e}}_{ji}$. To enable fast tensor product operations in large atomic structures, we adopt eSCN convolutions (Passaro & Zitnick, 2023), with incoming messages rotated to align with the $z$-axis during these operations. Non-linearity is then introduced through a gate activation layer. On top of this, we also include multi-headed attention mechanisms, as in EquiformerV2 (Liao et al., 2023), which allows our network to learn better from highly dense local atomic environments with varying node degree. Finally, the resulting output messages are rotated back to their original orientations, aggregated onto the node $i$, and passed through a feedforward network to update its output embedding ${\bm{n}}_{i}$ corresponding to onsite (diagonal) Hamiltonian blocks.
To learn offsite (non-diagonal) blocks, we use concepts from the Hamiltonian prediction networks in (Gong et al., 2023),(Wang et al., 2024a), and introduce learnable embeddings for every edge, defined between pairs of atoms located within a distance rcut from each other. The updated node embeddings are used to update the edges through a similar process, without the attention layer. The predicted outputs are then post-processed back into the uncoupled basis with the reverse transformation: $T^{-1}({\bm{H}}_{\gamma\delta}^{ij})$ = ${\bm{H}}_{\alpha\beta}^{ij}$, such that they can be used to reconstruct the Hamiltonian matrix block-by-block.
@@ 3.3 Training on augmented partitions
The combination of large graphs, dense connectivity, and large tensor representations of nodes and edges results in high memory consumption and long computation times per epoch during training (see Appendix H.2). If a distributed computing environment was used, the communication of intermediate values during the forward pass would incur significant overhead, impeding scalability (Wan et al., 2022). At the same time, graphs cannot be arbitrarily partitioned by removing inter-atomic connections. Doing so misinforms the network as it tries to fit the target data while aggregating inputs through an incorrect/incomplete graph structure. The inability to divide the graph into batches also slows down the training process overall.
@@ 3.4 Dataset creation
To generate sufficiently rich training data, existing datasets typically sample molecules at various time steps of molecular dynamics (MD) trajectories (Yu et al., 2023a; Schütt et al., 2019; Christensen & von Lilienfeld, 2020) or generate multiple small perturbations of the atoms in a crystalline lattice (Li et al., 2022). Here, as a representative subset of realistic (disordered) materials, we consider amorphous crystals and take advantage of the fact that (1) every node has a different local atomic environment, and (2) each structure contains a large sampling of different motifs. A wide range of training data can thus be captured within a single sample.
@@ 4.1 Ablation studies of the training approach
We study the model’s ability to generalize to different configurations of large systems by predicting the Hamiltonian matrix ${\bm{H}}$ of a full a-HfO2 sample (structure 3 in Table 1), which remains unseen during the training process. For all subsequent HfO2 experiments, the augmented partitioning scheme is only applied during training on structure 2, while the ${\bm{H}}$ of the full unseen structure is predicted during inference (structure 3). We use an rcut of 8 Å. Errors are reported separately for nodes ($\epsilon_{n}$) and edges ($\epsilon_{e}$) to distinguish intra- and inter-atomic orbital interactions, which typically have very different magnitudes.
@@
$n^{\prime}\;\;\;\;\;n$ 5.39 0.55
$n^{\prime}\rightarrow n$ 2.16 0.16
@@
First, we demonstrate the improvement in accuracy resulting from the augmented partitioning approach introduced in Section 3.3. In particular, we examine the influence of virtual nodes/edges in Table 2. Compared to training with raw partitions, the addition of virtual nodes and edges reduces the node ($\epsilon_{node}$) and edge prediction error ($\epsilon_{edge}$) by $\sim$60and $\sim$70respectively. Such an improvement is expected, as raw partitions omit a large proportion of boundary edges and thus incorrectly capture atomic neighborhoods. We note that as $t_{slice}<r_{cut}$, the one-hop neighborhood connects all nodes within each partition.
Next, we explore the impact of augmented partitioning on accuracy by training on an increasingly sub-divided graph in Table 3. The total number of labeled atoms used for training remains the same (3,000), while the total number of labeled edges decreases with more partitions. Despite the different divisions ranging from 5 ($t_{slice}\simeq$12 Å) to 27 ($t_{slice}\simeq$2 Å) slices, $\mathbf{\epsilon_{n}}$ and $\mathbf{\epsilon_{e}}$ remain very close to the values obtained by training with the full graph ($t_{slice}$ = 52.346 Å). The prediction error is thus insensitive to partition size. For small slices in this case, the reduced fraction of labeled connections along the $x$ direction does not affect the accuracy, as the remaining data along the $y$ and $z$ directions is sufficient to train the network.
@@ 4.2 Structural disorder
The strictly local architecture trained on augmented partitions performs consistently across the different test datasets. In all cases, prediction errors remain relatively insensitive to the partition size (Table 11 in Appendix I). Partitioning also allows us to extend $r_{cut}$ further (to 12 Å  for HfO2) while remaining within GPU memory limitations, effectively encompassing all non-zero matrix elements. As a result, the best achievable error for each material system ranges from 2.17 meV to 2.58 meV, for test structures containing 1000-3000 atoms and 200,000+ to 1,792,760+ edges. These values are comparable to what a previous study obtained (2.2 $meV$) using equivariant GNNs for smaller structures with $\leq$150 atoms per unit cell (Wang et al., 2024b).
@@ 4.3 Compositional disorder
To demonstrate our network’s ability to treat such problems, we train a model on a single sub-stoichiometric a-HfO1.8 structure where 10% of the oxygen atoms were replaced by randomly distributed vacancies represented as ghost atoms. We then use it to predict unseen full structures with a stoichiometry of a-HfO1.7, a-HfO1.8, and a-HfO1.9. The $\epsilon_{n}$ and $\epsilon_{e}$ (Table 5) remain within a small range (2.34-2.45 $mE_{h}$ and 0.15-0.16 $mE_{h}$, respectively), regardless of the vacancy concentration and distribution, showing that the network generalizes very well to compositional disorder, despite being trained on only one configuration. Models trained on other vacancy concentrations perform similarly. Their results, along with details of the vacancy implementation, can be found in Appendix I.1.